# ST-Think: How Multimodal Large Language Models Reason About 4D Worlds from Ego-Centric Videos

### v1

@@ 4.2 Post-training using GRPO
Revisit Group Relative Policy Optimization. Group Relative Policy Optimization(GRPO) is a recent policy optimization technique tailored for training large models (like language or multimodal models) with complex reasoning tasks. At its core, GRPO is similar in spirit to Proximal Policy Optimization (PPO) but introduces key differences in how the policy is updated and how the reward signal is used. Rather than relying on a learned value function (critic) to estimate expected reward, GRPO entirely foregoes the critic model. Instead, it estimates the baseline or expected reward by sampling a group of outputs from the current policy for each input and using their collective reward statistics. In other words, the model generates multiple candidate answers for a given query or state and uses their relative rewards to decide how to adjust the policy. This group-based advantage estimation removes the need to train a separate value network, simplifying the algorithm and reducing computational overhead. The development of ST-R1’s training procedure was inspired by the DeepSeek-R1 project, an advanced language reasoning model. DeepSeek-R1[10] demonstrated that reinforcement learning can be remarkably effective for honing a model’s reasoning abilities, even to the point of learning complex reasoning with minimal supervised data. It highlighted how GRPO can be used effectively in language model training to improve reasoning. The success of DeepSeek-R1 informed our approach for ST-R1: we similarly employ a two-stage training (CoT SFT followed by GRPO-based RL) to develop strong spatio-temporal reasoning capabilities.
@@
Reward Function Design. A critical aspect in this RL stage is the design of the reward function $R(x,y)$, especially for a complex domain like video understanding. We employ two complementary strategies to construct effective rewards, leveraging the question-answering (QA) format of the tasks.
Multiple-Choice QA Formulation: Many of the video understanding tasks are formulated as multiple-choice questions to simplify evaluation. Instead of having the model generate a free-form answer that might be hard to judge automatically, we provide a fixed set of answer options (one correct and the rest incorrect) for a given question. The model’s answer, in this case, is its choice among these options. The reward is then defined in a straightforward way: The model receives a positive reward if it selects the correct option and a zero (or negative) reward if it selects an incorrect option. The multiple-choice formulation thus greatly eases the reward engineering problem for many questions, allowing ST-R1 to learn from explicit right/wrong feedback.
AI-Generated Answer Scoring: Not all queries can be neatly converted into multiple-choice; some require descriptive answers or explanations. For such open-ended questions, we incorporate an AI-based scoring mechanism to assign rewards. We first obtain a reference answer or expected outcome for the question. We use a trained neural network to compare ST-R1’s answer $y$ with the reference. Using an AI judge or scoring function provides a flexible way to handle complex answers, essentially automating the reward assignment for cases where a simple rubric is not available.
Using the above reward strategies, we train ST-R1 with GRPO by continually generating answers and adjusting the policy. Over many iterations, this process increases the likelihood of the model producing answers that yield high rewards – in other words, answers that are correct and well-reasoned. Thanks to the multiple-choice design, whenever applicable, the model receives very clear signals about which answers are right. With the AI-based scoring, the model can also improve on free-form explanatory tasks by pleasing the automated judge. By combining these, we cover a wide spectrum of spatio-temporal reasoning tasks during RL training. The GRPO algorithm’s group-based updates ensure that the model explores different possible answers and learns from comparing them.
@@ 5.1 Evaluation Setup
Benchmark Models. We comprehensively evaluate 10 MLLMs from a diverse model family, including both closed-source and open-source models (the open-source models were tested in large and small-parameter versions, respectively). For proprietary models, we considered openai o3mini and Google Gemini-2.0 and Gemini-1.5 Pro. For open-source models, we evaluated models from InternVL2.5 [5], Qwen2-VL [30] , Qwen2.5-VL [1] and LLaVA-OneVision [13]. All evaluations were performed in a zero-shot setting and used the default cue for each model.
Metric Design. Two types of questions are included in our Ego-ST bench, including Open-ended Questions and Multiple-Choice Questions (see Figure 3). Regarding Multiple-Choice Question (MCQ), we follow standard workflow by using Accuracy ($\mathcal{ACC}$), based on exact matching, as the primary metric.
For the open-ended questions we used standard workflow [33, 21] by using by using LLM judge.We used GPT-4o api and set up a set of spatio-temporal evaluation of route description prompt.The evaluation method we devised for route descriptions is different from the evaluation methods, such as Video-ChatGPT [21], that use absolute ‘yes’ or ‘no’ and single-score evaluation methods. In total, we consider three aspects of route descriptions, as shown in Figure 7. Our approach includes direction changes, landmark changes, and logical semantic consistency. Each section is scored in integer from 0-5, and a final percentage score is calculated for the overall average score.
@@ 5.2 Results
Table 6 shows the overall performance of the model on the Ego-ST bench. Our main observations regarding the spatio-temporal reasoning capabilities of the different models are as follows:
Overall model performance. Overall, we have divided into two categories of models, which include Proprietary Models (using commercial api interfaces) and Open Source Models (using local deployments). In addition to evaluating multimodal large language models which support video, we have specifically selected models with strong reasoning capabilities for evaluation (e.g., o3-mini and Gemini-2.0). As shown in Table 6, we counted the overall scores and the scores of the strong spatio-temporal reasoning task (which includes route descriptions and direction change descriptions) for each model. We can see that under the same test environment and settings, all open source models do not score more than 60 in the strong spatio-temporal reasoning task, but perform very well in both the single spatial comprehension (landmark) and temporal comprehension (action). This reflects that the current multimodal large language modelling frameworks can effectively perceive spatial and temporal single tasks well respectively during training on large amounts of video and image data. From the human point of view, we are always in a 4D environment (3D+time), so the spatial and temporal reasoning and perception ability of the model is something we need to explore and improve in the future.
Mutiple Choice Question. We constructed three new types of multiple choice questions based on the open-ended question labels, as shown in Figure 3. The model needs to select the one with the correct chronological order and change among multiple options. These three types of problems cover direction change judgement, landmark scene change judgement and action change judgement. We consider these questions to be spatio-temporal reasoning questions in a wide sense, but each category has a different focus for evaluating the ability of the model.
@@
As shown in Table 6, It can be seen that the open-source model perform relatively poorly in direction change, which also indicates that the existing general multimodal large language model lacks of spatio-temporal reasoning ability, but performs relatively well in the lanmark and action, which is also due to the fact that the models nowadays have carried out in-depth research on both spatial and temporal understanding. But when these two task combine together the results are bad. In addition, based on the overall performance of each model in different tasks, we can see that the increase in the number of model parameters does not lead to a comprehensive improvement in spatio-temporal reasoning ability under the existing multimodal large language model training methods and frameworks. This is the reason why we would like to try to go for the improvement of spatio-temporal reasoning ability of multimodal large language models.
Model Best Part Avg. Direction (D) Landmark (L) Action (A)
Qwen2-VL-7B A 65.4 / 68.4 51.8 / 61.5 69.1 / 66.9 75.4 / 76.7
Qwen2-VL-7B-sft A 66.2 / 68.1 53.3 / 60.8 69.5 / 66.2 75.7 / 77.3
SFT+GRPO L 80.3 / 83.1 76.4 / 83.7 84.0 / 84.4 80.4 / 81.1
ST-CoT A 53.2 / 55.3 47.6 / 48.3 52.5 / 56.1 59.5 / 61.5
ST-R1 A 86.3 / 86.1 82.0 / 85.5 88.1 / 85.5 88.9 / 87.2
@@
Open-ended Route Description. As illustrated in Table 1, we present detailed scores for each component within the comprehensive route description task. It is evident that the scores for the direction and landmark categories in the open-ended Route Description task are substantially lower compared to their corresponding scores in the multiple-choice question tasks. By further comparing the differences in performance between the complete route description and three distinct multiple-choice conditions (direction, landmark, and action), we observed that the complete description task poses significantly greater difficulty than the single-dimensional multiple-choice tasks. This disparity highlights that generating complete route descriptions necessitates a more sophisticated integration of overall spatio-temporal information, deeper reasoning regarding temporal causality, and flexible shifts in spatial perspective. Conversely, the multiple-choice tasks predominantly assess recognition and recall of isolated spatio-temporal details. Consequently, the complete route description task, structured as an open-ended question-answering scenario, presents unique challenges that demand higher-order spatial-temporal reasoning capabilities from models.
Furthermore, model performance in the open-ended description task generally aligns with their performance on the multiple-choice questions. Specifically, across evaluated open-source models, scores in the Landmark category under the Spatial Semantic Understanding dimension are consistently the highest. In contrast, performance in the Direction Changing Environment, representative of Spatio-Temporal Reasoning, remains relatively poor among both open-source and closed-source models. This result aligns with observations from earlier multiple-choice evaluations, underscoring that current multimodal large language model architectures and associated training paradigms exhibit notable shortcomings in effectively handling complex spatio-temporal reasoning tasks.
@@
5.3 Results of our ST-R1 Model
Based on our thinking about the benchmark evaluation results and Deepseek-R1 training paradigm, we propose a new training paradigm for spatio-temporal reasoning in video. We divide the multiple-choice questions in the overall benchmark into Train and Test in the ratio of 1:6. we aim to explore whether the reasoning training framework combining spatio-temporal CoT fine-tuning and GRPO reinforcement learning has stronger generalisation and spatio-temporal reasoning ability with a small amount of training data.
As shown in Table 2, the two-stage training approach for the ST-R1 model yields superior generalization performance even with limited training data. Our ST-R1 model, which is post-trained based on Qwen2-VL-7B, achieves significant performance improvements over traditional SFT methods. Initially, the model is trained using spatio-temporal chain-of-thought (CoT) SFT as a cold start, followed by a secondary training phase employing GRPO. Although direct GRPO tuning improves performance, pre-training with CoT SFT further enhances the results. Experimental results also indicate that when using GRPO for reinforcement alignment, it is crucial to perform prompt-tuning with CoT SFT rather than with conventional SFT. To demonstrate the necessity of spatio-temporal CoT SFT, we also trained a model using regular SFT with GRPO without CoT, as shown in Figure 8. Under the same training conditions, the model trained with regular SFT+GRPO exhibited extremely unstable performance during GRPO training, and its final performance was markedly inferior to that achieved with the CoT SFT+GRPO configuration.
@@ 6 Conclusion and Future Work
In this paper, we explored the spatio-temporal reasoning capabilities of current multimodal large language models, highlighting this as a challenging yet essential research direction. To facilitate comprehensive evaluation and benchmarking, we introduced ST-Bench, a new dataset and benchmark designed specifically to assess existing models’ performance in complex spatio-temporal reasoning tasks. Our results indicate that the performance gap between open-source and closed-source models is minimal. However, we observed a significant performance drop when models are required to organize and generate integrated, open-ended route descriptions, despite their relatively high performance on isolated tasks such as recognizing landmarks or directional changes in multiple-choice settings. This underscores the critical challenge that current multimodal models still lack the ability to effectively integrate complex spatio-temporal information. Furthermore, we examined optimal post-training strategies for enhancing video-based spatio-temporal reasoning capabilities. Our proposed two-stage post-training approach demonstrated substantial improvements over existing methods, providing valuable insights not only for spatio-temporal reasoning but also for other video reasoning tasks.
In the future, several promising directions merit exploration. These include: (1) enriching multimodal large language models by incorporating additional modalities and diverse data sources to enhance spatio-temporal inference; (2) guiding models toward implicit reconstruction tasks, encouraging a deeper understanding of temporal and spatial dynamics; (3) efficiently developing generalized video reasoning models with robust spatio-temporal reasoning capabilities; and (4) advancing computationally efficient inference methods to support scalable deployment. Addressing these challenges will be critical to advancing the state-of-the-art in multimodal spatio-temporal reasoning.

### v2

@@ 4.2 Post-training using GRPO
Revisit Group Relative Policy Optimization. Group Relative Policy Optimization(GRPO) is a recent policy optimization technique tailored for training large models (like language or multimodal models) with complex reasoning tasks. At its core, GRPO is similar in spirit to Proximal Policy Optimization (PPO) but introduces key differences in how the policy is updated and how the reward signal is used. Rather than relying on a learned value function (critic) to estimate expected reward, GRPO entirely foregoes the critic model. Instead, it estimates the baseline or expected reward by sampling a group of outputs from the current policy for each input and using their collective reward statistics. In other words, the model generates multiple candidate answers for a given query or state and uses their relative rewards to decide how to adjust the policy. This group-based advantage estimation removes the need to train a separate value network, simplifying the algorithm and reducing computational overhead. The development of ST-R1’s training procedure was inspired by the DeepSeek-R1 project, an advanced language reasoning model. DeepSeek-R1[9] demonstrated that reinforcement learning can be remarkably effective for honing a model’s reasoning abilities, even to the point of learning complex reasoning with minimal supervised data. It highlighted how GRPO can be used effectively in language model training to improve reasoning. The success of DeepSeek-R1 informed our approach for ST-R1: we similarly employ a two-stage training (CoT SFT followed by GRPO-based RL) to develop strong spatial-temporal reasoning capabilities.
@@
Reward Function Design. A critical aspect in this RL stage is the design of the reward function $R(x,y)$, especially for a complex domain like video understanding. Many of the video understanding tasks are formulated as multiple-choice questions to simplify evaluation. Instead of having the model generate a free-form answer that might be hard to judge automatically, we provide a fixed set of answer options (one correct and the rest incorrect) for a given question. The model’s answer, in this case, is its choice among these options. The reward is then defined in a straightforward way: The model receives a positive reward if it selects the correct option and a zero (or negative) reward if it selects an incorrect option. The multiple-choice formulation thus greatly eases the reward engineering problem for many questions, allowing ST-R1 to learn from explicit right/wrong feedback.
Using the above reward strategies, we train ST-R1 with GRPO by continually generating answers and adjusting the policy. Over many iterations, this process increases the likelihood of the model producing answers that yield high rewards – in other words, answers that are correct and well-reasoned. Due to the multiple-choice reward design, the model receives very clear signals about which answers are right. The GRPO algorithm’s group-based updates ensure that the model explores different possible answers and learns from comparing them.
@@ 5.1 Evaluation Setup
Benchmark Models. We comprehensively evaluate 10 MLLMs from a diverse model family, including both closed-source and open-source models (the open-source models were tested in large and small-parameter versions, respectively). For proprietary models, we considered OpenAI-o3mini, Google Gemini-2.0, and Gemini-1.5 Pro. For open-source models, we evaluated models from InternVL2.5 [5], Qwen2-VL [27] , Qwen2.5-VL [1] and LLaVA-OneVision [11]. All evaluations were performed in a zero-shot setting and used the default cue for each model. More details of the parameters can be found in the Appendix.
Metric Design. Two types of questions are included in our Ego-ST bench, including Open-ended Questions and Multiple-Choice Questions (see Figure 3). Regarding Multiple-Choice Questions (MCQ), we follow standard workflow by using Accuracy ($\mathcal{ACC}$), based on exact matching, as the primary metric.
For the open-ended questions we used standard workflow [30, 19] by using by using LLM judge.We used GPT-4o api and set up a set of spatial-temporal evaluation of route description prompt.The evaluation method we devised for route descriptions is different from the evaluation methods, such as Video-ChatGPT [19], that use absolute ‘yes’ or ‘no’ and single-score evaluation methods. In total, we consider three aspects of route descriptions, as shown in Figure 7. Our approach includes direction changes, landmark changes, and logical semantic consistency. Each section is scored in integer from 0-5, and a final percentage score is calculated for the overall average score.
@@ 5.2 Results
Table 6 shows the overall performance of the model on the Ego-ST bench. Our main observations regarding the spatial-temporal reasoning capabilities of the different models are as follows:
Overall model performance. Overall, we have divided into two categories of models, which include Proprietary Models (using commercial api interfaces) and Open Source Models (using local deployments). In addition to evaluating multimodal large language models which support video, we have specifically selected models with strong reasoning capabilities for evaluation (e.g., o3-mini and Gemini-2.0). As shown in Table 6, we counted the overall scores and the scores of the strong spatial-temporal reasoning task (which includes route descriptions and direction change descriptions) for each model. We can see that under the same test environment and settings, all open source models do not score more than 60 in the strong spatial-temporal reasoning task, but perform very well in both the single spatial comprehension (landmark) and temporal comprehension (action). This reflects that the current multimodal large language modelling frameworks can effectively perceive spatial and temporal single tasks well respectively during training on large amounts of video and image data. From the human point of view, we are always in a 4D environment (3D+time), so the spatial and temporal reasoning and perception ability of the model is something we need to explore and improve in the future.
Mutiple Choice Question. We constructed three new types of multiple choice questions based on the open-ended question labels, as shown in Figure 3. The model needs to select the one with the correct chronological order and change among multiple options. These three types of problems cover direction change judgement, landmark scene change judgement and action change judgement. We consider these questions to be spatial-temporal reasoning questions in a wide sense, but each category has a different focus for evaluating the ability of the model.
@@
As shown in Table 6, It can be seen that the open-source model perform relatively poorly in direction change, which also indicates that the existing general multimodal large language model lacks of spatial-temporal reasoning ability, but performs relatively well in the lanmark and action, which is also due to the fact that the models nowadays have carried out in-depth research on both spatial and temporal understanding. But when these two task combine together the results are bad. In addition, based on the overall performance of each model in different tasks, we can see that the increase in the number of model parameters does not lead to a comprehensive improvement in spatial-temporal reasoning ability under the existing multimodal large language model training methods and frameworks. This is the reason why we would like to try to go for the improvement of spatial-temporal reasoning ability of multimodal large language models.
Model Training Num. Avg. Direction (D) Landmark (L) Action (A)
Qwen2-VL-7B 0/3582 65.4 / 68.4 51.8 / 61.5 69.1 / 66.9 75.4 / 76.7
Qwen2-VL-7B-sft 630/3582 66.2 / 68.1 53.3 / 60.8 69.5 / 66.2 75.7 / 77.3
SFT+GRPO 630/3582 80.3 / 83.1 76.4 / 83.7 84.0 / 84.4 80.4 / 81.1
ST-CoT 630/3582 53.2 / 55.3 47.6 / 48.3 52.5 / 56.1 59.5 / 61.5
ST-R1 63/3582 64.8 / 63.5 54.9 / 53.8 70.9 / 71.0 68.6 / 65.6
ST-R1 126/3582 62.4 / 62.1 53.0 / 63.5 63.4 / 63.0 70.7 / 59.8
ST-R1 252/3582 67.1 / 66.2 59.8 / 68.2 67.5 / 66.9 73.9 / 63.6
ST-R1 378/3582 70.0 / 67.8 56.4 / 51.8 80.4 / 82.3 73.3 / 69.4
ST-R1 504/3582 80.7 / 82.6 75.3 / 83.2 86.5 / 84.1 80.4 / 80.5
ST-R1 630/3582 86.3 / 86.1 82.0 / 85.5 88.1 / 85.5 88.9 / 87.2
@@
Open-ended Route Description. As illustrated in Table 1, we present detailed scores for each component within the comprehensive route description task. It is evident that the scores for the direction and landmark categories in the open-ended Route Description task are substantially lower compared to their corresponding scores in the multiple-choice question tasks. By further comparing the differences in performance between the complete route description and three distinct multiple-choice conditions (direction, landmark, and action), we observed that the complete description task poses significantly greater difficulty than the single-dimensional multiple-choice tasks. This disparity highlights that generating complete route descriptions necessitates a more sophisticated integration of overall spatial-temporal information, deeper reasoning regarding temporal causality, and flexible shifts in spatial perspective. Conversely, the multiple-choice tasks predominantly assess recognition and recall of isolated spatial-temporal details. Consequently, the complete route description task, structured as an open-ended question-answering scenario, presents unique challenges that demand higher-order spatial-temporal reasoning capabilities from models.
Furthermore, model performance in the open-ended description task generally aligns with their performance on the multiple-choice questions. Specifically, across evaluated open-source models, scores in the Landmark category under the Spatial Semantic Understanding dimension are consistently the highest. In contrast, performance in the Direction Changing Environment, representative of spatial-Temporal Reasoning, remains relatively poor among both open-source and closed-source models. This result aligns with observations from earlier multiple-choice evaluations, underscoring that current multimodal large language model architectures and associated training paradigms exhibit notable shortcomings in effectively handling complex spatial-temporal reasoning tasks.
@@
5.3 Results of ST-R1 Model
Comparison of different training paradigms. Based on our thinking about the benchmark evaluation results and Deepseek-R1 training paradigm, we propose a new training paradigm for spatial-temporal reasoning in video. We divide the multiple-choice questions in the overall benchmark into Train and Test in the ratio of 1:6. In addition, we incrementally expanded the training dataset from small to large, aiming to investigate whether our reasoning framework—combining spatial-temporal Chain-of-Thought (CoT) fine-tuning with GRPO reinforcement learning—can achieve superior generalization and spatial-temporal reasoning capabilities, even with limited training data.
As shown in Table 2, the two-stage training approach for the ST-R1 model yields superior generalization performance even with limited training data. Our ST-R1 model, which is post-trained based on Qwen2-VL-7B, achieves significant performance improvements over traditional SFT methods. Initially, the model is trained using spatial-temporal chain-of-thought (CoT) SFT as a cold start, followed by a secondary training phase employing GRPO. Although direct GRPO tuning improves performance, pre-training with CoT SFT further enhances the results. Experimental results also indicate that when using GRPO for reinforcement alignment, it is crucial to perform prompt-tuning with CoT SFT rather than with conventional SFT. To demonstrate the necessity of spatial-temporal CoT SFT, we also trained a model using regular SFT with GRPO without CoT, as shown in Figure 8. Under the same training conditions, the model trained with regular SFT+GRPO exhibited extremely unstable performance during GRPO training, and its final performance was markedly inferior to that achieved with the CoT SFT+GRPO configuration.
Out of Distribution (OOD) Generalization. The experiments in the previous sections showed that a small amount of training data can generalise strongly in same distribution. As shown in Table 3, we continue to experiment with a small amount of training data (630 samples) to test on OOD, and we also validate the performance of indoor to outdoor across domains. As we can see from the experimental results, ST-R1 is able to perform well on OOD with a small amount of training data, and ST-R1 achieves a significant improvement of over 32% in the two types of OOD tasks, respectively. Through the comprehensive analysis of Table 2 and Table 3, we can see that ST-R1 is able to show an extremely strong and comprehensive generalisation ability with a small amount of training data.
Model Train Test Direction (D) Landmark (L) Avg.
Qwen2-VL-7B - SUN3D 45.0 / 49.7 73.0 / 73.8 59.0 / 61.8
ST-R1(indoor-cross) HUJI+Self+Aria SUN3D 74.8 / 90.1 83.5 / 80.4 79.2 / 85.3
Qwen2-VL-7B - HUJI+Self 56.6 / 77.0 63.9 / 73.9 60.3 / 75.5
ST-R1(indoor-outdoor) SUN3D+Aria HUJI+Self 77.6 / 100 81.5 / 84.1 79.6 / 92.1
Table 3: OOD Testing. We have done detailed tests on indoor scenarios OOD data and indoor to outdoor performance migration and generalization.
@@ 6 Conclusion And Future Work
In this paper, we explored the spatial-temporal reasoning capabilities of current multimodal large language models, highlighting this as a challenging yet essential research direction. To facilitate comprehensive evaluation and benchmarking, we introduced ST-Bench, a new dataset and benchmark designed specifically to assess existing models’ performance in complex spatial-temporal reasoning tasks. Our results indicate that the performance gap between open-source and closed-source models is minimal. However, we observed a significant performance drop when models are required to organize and generate integrated, open-ended route descriptions, despite their relatively high performance on isolated tasks such as recognizing landmarks or directional changes in multiple-choice settings. This underscores the critical challenge that current multimodal models still lack the ability to effectively integrate complex spatial-temporal information. Furthermore, we examined optimal post-training strategies for enhancing video-based spatial-temporal reasoning capabilities. Our proposed two-stage post-training approach demonstrated substantial improvements over existing methods, providing valuable insights not only for spatial-temporal reasoning but also for other video reasoning tasks.
Moving forward, several promising directions merit exploration. These include: (1) enriching multimodal large language models by incorporating additional modalities and diverse data sources to enhance spatial-temporal inference; (2) guiding models toward implicit reconstruction tasks, encouraging a deeper understanding of temporal and spatial dynamics; (3) efficiently developing generalized video reasoning models with robust spatial-temporal reasoning capabilities; and (4) advancing computationally efficient inference methods to support scalable deployment.
Addressing these challenges will be critical to advancing the state-of-the-art in multimodal spatial-temporal reasoning.